# Segmented Operations using Matrix Multiplications

### v1

@@ 3.2 The balance between the MMU and VCU
One such operation that “troubles” the MMU is a simple block-reordering of a large vector. Assume that we have a vector of size $ms=ns^{2}$, which is partitioned in $ns$ blocks of size $s$, and the blocks are grouped in consecutive groups of $s$ blocks each. The goal is to reverse the order of the blocks within each group.
Assume that we only have an MMU available. The rows of a matrix can only be permuted with multiplications from the left. From the definition of the MMU, we can reverse the order of at most $s$ blocks (a single group) at a time with a single matrix multiplication. This means that we need $\Omega(n)$ total calls to the MMU. On the other hand, this reordering can be easily implemented with an AC[0] circuit. ∎
3.3 Circuit analysis of the vector unit
In MMV-RAM, the VCU can execute a fixed set of operations that belong to AC[0], in particular, uniform families of unbounded fan-in circuits with constant depth and polynomial size in the input vector length, $n$. In Table 3.3, we summarize all the vector instructions that are used in Algorithm 1. All operations involve one or two vectors of size $n$ with single-bit elements or $B$-bit integers. Instructions $6$-$9$ are parametrized by $s$, the (fixed) model parameter. The instructions can be implemented with uniform AC[0] circuits for every input size $n$ as follows.
  • Element-wise $\mathsf{ADD}/\mathsf{SUB}$ can also be achieved with AC[0] circuits, with $\mathcal{O}(nB^{2})$ and $\mathcal{O}(B)$ fan-in boolean gates. See e.g. [57, Chapter 1] or [20, Chapter 8] for further details and for more advanced circuits and techniques. We note that, in the main algorithm, we only use subtractions $a-b$ for the special case where $a\geq b$.
  • $\mathsf{SCATTERS}/\mathsf{GATHERS}$ are also parametrized by $s$. As the name suggests, $\mathsf{GATHERS}$ copies the last element of the $i$-th $s$-segment of $\bm{\mathrm{x}}$, to $\bm{\mathrm{z}}(i-1)$. Accordingly, $\mathsf{SCATTERS}$ copies the $(i-1)$-th element of $\bm{\mathrm{x}}$ to the last position of the $i$-th $s$-segment of $\bm{\mathrm{z}}$. The corresponding circuits have size $\mathcal{O}(sB)$, since they only copy one element per segment to a single position in the output, and the fan-in/-out of the gates are all constant.
  • Finally, the instruction $\mathsf{REVSPEC}$ is a uniform circuit that corrects the misspeculated (unsegmented) scans in Algorithm 1. This circuit is described in detail in Section 4, specifically, in Lemma 4.1.
4 Segmented scan in MMV-RAM
This section presents a parallel segmented scan in the MMV-RAM model, Algorithm 1. The algorithm follows a block-recursive approach described next, with the key feature of exclusively employing vector operations and matrix multiplications.
4.1 A block recursive approach to segmented scan
@@
Here, we outline a block-recursive approach to computing segmented scans, which is the basis of Algorithm 1. The core idea of computing segmented scans (recursively) via a contraction argument is well known in the literature; see [5] and [49, Algorithms 3 and 4] for fan-in of size two. Notably, for the unsegmented scan problem, [64] extended the contraction approach to blocks (or fan-in) of size $s$, generalizing the classic Brent–Kung algorithm [11]. In the following, we briefly describe a similar generalization for the segmented case.
@@
Given an input vector $\bm{\mathrm{x}}$ and a boolean vector $\bm{\mathrm{f}}$ both of size $n$, the block recursive approach initially partitions $\bm{\mathrm{x}}$ and $\bm{\mathrm{f}}$ in consecutive blocks of size $s$. Then the following steps are executed:
@@
    On each $s$-block of $(\bm{\mathrm{x}},\bm{\mathrm{f}})$, compute (in parallel) the “local” segmented scan and store the result in place on $\bm{\mathrm{x}}$.
@@
  3. 3.
    Generate a vector $\bm{\mathrm{f}}_{s}$ of size $\lceil\tfrac{n}{s}\rceil$ by applying the logical OR operator to the elements of each $s$-block of $\bm{\mathrm{f}}$.

### v2

@@ 3.2 The balance between the MMU and VCU
Having expressed the basic operations as matrix-vector products, we can now derive lower bounds. Let us study how to perform the following block-reordering operation of a large vector. Assume a vector $\bm{\mathrm{x}}$ of size $n=ms^{2}$, for some $m$, which is partitioned in $ms$ blocks of size $s$ each, and the blocks are grouped in consecutive groups of $s$ blocks each. The goal is to reverse the order of the blocks within each group. Algebraically, this operation can be written as follows:
$\bm{\mathrm{y}}\leftarrow\left(\bm{\mathrm{I}}_{m}\otimes\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s}\right)\bm{\mathrm{x}},$
where $\bm{\mathrm{J}}_{s}$ is the identity matrix with its columns in reverse order. Note that this reordering can easily be implemented with an AC[0] circuit.
Assume that we have an MMV-RAM algorithm which performs this reordering using only the aforementioned operations. Based on the above, this algorithm can be written as a sequence of matrix-vector products
$\bm{\mathrm{y}}\leftarrow\bm{\mathrm{L}}_{1}\bm{\mathrm{R}}_{1}\bm{\mathrm{L}}_{2}\bm{\mathrm{R}}_{2}\ldots\bm{\mathrm{L}}_{k}\bm{\mathrm{R}}_{k}\bm{\mathrm{x}},$
where $\bm{\mathrm{L}}_{i}$ and $\bm{\mathrm{R}}_{i}$ are left and right multiplication operations as defined above (some of them might be equal to the identity, in which case these operations are ignored). Now we will derive an $\Omega(s)$ lower bound on the number of operations that need to be executed.
First, note that the equality holds necessarily:
$\bm{\mathrm{L}}_{1}\bm{\mathrm{R}}_{1}\bm{\mathrm{L}}_{2}\bm{\mathrm{R}}_{2}\ldots\bm{\mathrm{L}}_{k}\bm{\mathrm{R}}_{k}=\bm{\mathrm{I}}_{m}\otimes\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s},$
because the transformation must return the correct result for every input vector. Therefore, it suffices to argue about the minimum number of $\bm{\mathrm{L}}$ and $\bm{\mathrm{R}}$ factors required to produce the matrix $\bm{\mathrm{I}}_{m}\otimes\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s}$.
We can now observe that a sequence of transformations $\bm{\mathrm{R}}_{1}\bm{\mathrm{R}}_{2}\ldots\bm{\mathrm{R}}_{k}$ cannot produce the matrix $\bm{\mathrm{I}}_{m}\otimes\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s}$ with less than $k=\Omega(s)$ products. This is because the matrices $\bm{\mathrm{R}}_{i}$ are banded with bandwidth s1s-1, therefore, each product increases the bandwidth by at most $s$. However, the matrix $\bm{\mathrm{I}}_{m}\otimes\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s}$ has a bandwidth of $s^{2}$.
Next, note that, due to its structure, a transformation $\bm{\mathrm{L}}_{i}$ operates on at most $s^{2}$ consecutive elements of $\bm{\mathrm{x}}$. Assume now that we have executed a sequence of transformations $\bm{\mathrm{L}}_{1}\bm{\mathrm{R}}_{1}\bm{\mathrm{L}}_{2}\bm{\mathrm{R}}_{2}\ldots\bm{\mathrm{L}}_{k}\bm{\mathrm{R}}_{k}\bm{\mathrm{x}}$. Based on the previous argument, there are at least $n-ks^{2}$ elements of $\bm{\mathrm{x}}$ that have only been modified by the $\bm{\mathrm{R}}_{i}$ transformations. If $n$ is large enough, this means that there is at least an entire block $\bm{\mathrm{J}}_{s}\otimes\bm{\mathrm{I}}_{s}$ which has has not been modified by $\bm{\mathrm{L}}_{i}$ transformations. Based on the argument above, we need $k=\Omega(s)$ left and right transformations to reorder this block, concluding the proof. ∎
Combining all these arguments together, we can state the following theorem.
Theorem 3.5.
In MMV-RAM, the MMU and the VCU complement each other in the following sense:
  • The VCU can compute non-linear element-wise functions, such as RELU, on a vector $\bm{\mathrm{x}}$ in a single step, while the MMU cannot execute such operations.
  • The MMU can multiply two matrices $\bm{\mathrm{A}}$ and $\bm{\mathrm{B}}$ with sizes $n\times s$ and $s\times s$, in one step. However, any MMV-RAM algorithm that uses only the scalar and vector units to compute $\bm{\mathrm{A}}\bm{\mathrm{B}}$, and executes work that scales polynomially in $s$, requires $\Omega\left(\frac{\log(s)}{\log(\log(s))}\right)$ steps.
  • There exists a linear vector transformation (a permutation) which can be efficiently implemented in the VCU in one step and linear work, but it requires $\Omega(s)$ calls to the MMU.
3.3 Extensions of MMV-RAM
Here we briefly discuss possible variants and extensions of the MMV-RAM model.
3.3.1 Different circuit classes for the VCU
The VCU of MMV-RAM consists of AC[0] circuits. Other circuit complexity classes might be worth investigating, and what is their effect in the analysis of algorithms. As an example, recall that our main algorithms for segmented operations rely on gates with fan-in and fan-out of $\mathcal{O}(\operatorname{poly}\left(s,B\right))$ size, i.e., independent of input size. Some hardware implementations, however, might only support constant fan-in gates. As such, NC[0] can be investigated as an alternative for the VCU. Other classes can also be considered depending on the applications. Some other examples from [61] include semi-bounded fan-in circuits (SAC), AC “with counters” (ACC), or AC[0] circuits extended with additional modulo gates. These classes lie between NC[0] and NC[1] in the so-called NC-hierarchy (cf. [61, Chapter 4]). In all cases, the complexity of VCU instructions (size-depth) and the balance with the MMU should be carefully calibrated.
3.3.2 Extending the right-hand-side of the MMU
$\begin{pmatrix}\bm{\mathrm{A}}_{1}\bm{\mathrm{B}}_{1}\\ \vdots\\ \bm{\mathrm{A}}_{k}\bm{\mathrm{B}}_{k}\end{pmatrix}\leftarrow\begin{pmatrix}\bm{\mathrm{A}}_{1}\\ \vdots\\ \bm{\mathrm{A}}_{k}\end{pmatrix}\odot\begin{pmatrix}\bm{\mathrm{B}}_{1}\\ \vdots\\ \bm{\mathrm{B}}_{k}\end{pmatrix}$,
instead of
$\begin{pmatrix}\bm{\mathrm{A}}_{1}\bm{\mathrm{B}}\\ \vdots\\ \bm{\mathrm{A}}_{k}\bm{\mathrm{B}}\end{pmatrix}\leftarrow\begin{pmatrix}\bm{\mathrm{A}}_{1}\\ \vdots\\ \bm{\mathrm{A}}_{k}\end{pmatrix}\bm{\mathrm{B}}.$
4 Segmented operations in MMV-RAM
Given the definition of MMV-RAM, we next proceed to algorithmic design for segmented operations and their applications. Our main algorithms rely on unsegmented scan as a parallel primitive to perform speculative computations. For this, we use the unsegmented scan algorithm of [69], which is a generalization of the Brent–Kung algorithm [11]. Theorem 4.1 reports the complexity of this algorithm in the MMV-RAM model. The work-depth bounds match the ones of [69] with respect to $n$ and $s$, but additionally here the MMV-RAM model takes into account the gather/scatter vector operations explicitly.
Theorem 4.1 (Scan).
Given a vector $\bm{\mathrm{x}}$ with integer entries bounded by $\bm{\mathrm{x}}_{i}\leq M$, for some $M$, we can compute a vector $\bm{\mathrm{z}}$ that contains the scan of $\bm{\mathrm{x}}$ in the MMV-RAM model using the algorithm of [69] (listed in Algorithm 1). It requires:
$\mathcal{O}(\log_{s}(n))\text{ steps, }\mathcal{O}\left(\mathcal{M}(\tfrac{n}{s})+nB^{2}\right)\text{ work},$
and $B\in\mathcal{O}(\log(nM))$ bits-per-element to avoid overflow. Any algorithm that uses only the VCU and executes $n^{\mathcal{O}(1)}$ work requires $\Omega\left(\frac{\log(n)}{\log(\log(n))}\right)$ steps.
Proof.
$W(t)$ $=W(t+1)+2\mathcal{M}(\tfrac{n}{s^{t}})+\mathcal{O}(sB)$
$=\sum_{t=1}^{\lfloor\log_{s}(n)\rfloor}\mathcal{M}(\tfrac{n}{s^{t}})+\mathcal{O}(sB\log_{s}(n))$
$\in\mathcal{O}\left(\mathcal{M}(\tfrac{n}{s})+sB\log_{s}(n)\right).$
The total number of bits used to store numbers during the execution of the algorithm is bounded as follows. In iteration $t=1,\ldots,\log_{s}(n)$, we have two $\mathsf{MATMUL}$ operations. These are the only operations that can increase the magnitude of the elements. Let $M(1)=M$, and denote by $M(t)$ the magnitude of the largest element at step $t$. Since the right-hand-side matrices have only zeros and ones, then every matrix multiplication can increase the magnitude of the input elements at most by a factor of $s$. Thus:
which means that $\mathcal{O}(\log(n)+\log(M))$ bits are sufficient to avoid overflow. ∎
1:procedure Scan($\bm{\mathrm{x}}$; $s$) $\triangleright$ $n\leftarrow\operatorname{len}(\bm{\mathrm{x}})$
2:  $\bm{\mathrm{z}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$
3:  if $n\leq s$ then
4:    return $\bm{\mathrm{z}}$
5:  $\bm{\mathrm{x}}_{s}\leftarrow\mathsf{GATHERS}(\bm{\mathrm{z}};s)$
6:  $\bm{\mathrm{z}}_{s}\leftarrow\textsc{Scan}(\bm{\mathrm{x}}_{s};s)$
7:  $\bm{\mathrm{z}}\leftarrow\mathsf{SCATTERS}(\bm{\mathrm{z}}_{s};s)$
8:  $\bm{\mathrm{z}}(s-1:n)\leftarrow\mathsf{MATMUL}(\bm{\mathrm{z}}(s-1:n),\bm{\mathrm{B}}_{s})$ $\triangleright$ $\bm{\mathrm{B}}_{s}$ is the identity with all-ones in the first row
9:  return $\bm{\mathrm{z}}$
Algorithm 1 Scan in MMV-RAM (simplified version of [69, Alg. 2])
4.1 Segmented Sum (SCD) and Segmented Scan (SSCR)
We next analyze our main algorithms for segmented operations. First, we present a segmented sum algorithm, called SCD (Scan, Compress, and Differentiation), specialized for matrix multiplication accelerators (Algorithm˜2). In the first step of SCD, the unsegmented scan of the input $\bm{\mathrm{x}}$ is computed speculatively. In the second step, $\bm{\mathrm{f}}$ is slightly modified to $\bm{\mathrm{f}}^{-}$ so that the end positions of all segments are marked by the ones in $\bm{\mathrm{f}}^{-}$. Indeed, $\bm{\mathrm{f}}^{-}$ is obtained by $\bm{\mathrm{f}}$ by shifting the vector to the left and appending 1 at the end. Then, the scanned values of $\bm{\mathrm{x}}$ corresponding to the end position of all segments are collected through a parallel compaction / compress operation using $\bm{\mathrm{f}}^{-}$. One can verify that the segmented sum can be obtained by applying vector differentiation $\bm{\mathrm{z}}_{\rm diff}(i)=\bm{\mathrm{z}}(i)-\bm{\mathrm{z}}(i-1)$ (assume $\bm{\mathrm{z}}(-1)=0$) on the output vector of compress.
Now, we argue that all three steps of SCD could benefit from a MMU, i.e., SCD is amenable to an efficient implementation in the MMV-RAM model. SCAN can be implemented using matrix multiplications as shown in [17, 69]. Similarly, parallel compress is typically implemented by first scanning the input flag vector to calculate the output indices. Finally, the DIFF step of SCD could also employ the MMU as as follows (albeit with low utilization of the MMU). Let $\bm{\mathrm{A}}$ be a row-wise matrix view with $s$ columns of an arbitrary vector $\bm{\mathrm{x}}$ (padded accordingly). Let
$\bm{\mathrm{D}}_{s}:=\begin{pmatrix}1&-1&0&\dots&0\\ 0&1&-1&\dots&0\\ 0&0&1&\dots&0\\ \vdots&0&0&\ddots&-1&\\ 0&0&\dots&0&1\end{pmatrix}_{s\times s}.$ (3)
It is easy to verify that the matrix product $\bm{\mathrm{A}}\bm{\mathrm{D}}_{s}$ computes the difference $\bm{\mathrm{x}}(i)-\bm{\mathrm{x}}(i-1)$ on each $s$-block except the boundary entries between blocks. The remaining entries on the boundary can be computed using the VCU.
To the best of our knowledge, the derivation of segmented sum as a composition of scan, compress, and vector differentiation, has not been explicitly discussed in the literature, but similar ideas exist. One relevant example is the “fast segmented sum”, Algorithm 6 of [40], which we call FSS from now on. Similarly to SCD, FSS performs an unsegmented scan on $\bm{\mathrm{x}}$. FSS requires an additional copy of the input data vector to a temporary vector. Next, using the additional temporary vector, FSS computes the segmented sums using a gather in a parallel for-loop. In contrast, and crucially, SCD does not require an additional temporary vector and could take advantage of the MMU for the index calculations of the parallel compaction/compress operation.
1:procedure SCD($\bm{\mathrm{x}}$, $\bm{\mathrm{f}}$)
[…]
3:  $\bm{\mathrm{f}}^{-}\leftarrow\bm{\mathrm{f}}(1:)$ appended with 11
5:  $\bm{\mathrm{z}}_{\rm diff}\leftarrow\text{DIFF}(\bm{\mathrm{z}})$ $\triangleright$ Vector Differentiation.
6:  return $\bm{\mathrm{z}}_{\rm diff}$
Algorithm 2 Segmented Sum $\cong$
DIFF $\circ$ COMPRESS $\circ$ SCAN
The Segmented Scan Algorithm˜3, dubbed SSCR (Scan, Scan, Compress, and Revert), shares a similar approach to the SCD algorithm. First, it performs a speculative scan of $\bm{\mathrm{x}}$ and then a parallel compress to collect the last entries of each segment. In addition, SSCR requires a scan of the boolean flag vector along with a parallel correction operation that could be performed in the VCU (Line 5 of Algorithm˜3). We call the vector-only correction operator REVERT, which takes as input $\widehat{\bm{\mathrm{x}}},\widehat{\bm{\mathrm{f}}}$, and $\bm{\mathrm{w}}$, and it subtracts from each segment the last element of the previous segment. A simple (parallel) implementation of REVERT is the following:
procedure REVERT($\widehat{\bm{\mathrm{x}}}$, $\widehat{\bm{\mathrm{f}}}$, $\bm{\mathrm{w}}$)
  parfor idx in $\text{len}(\bm{\mathrm{w}})$ do
   $\bm{\mathrm{z}}\leftarrow\widehat{\bm{\mathrm{x}}}-(\widehat{\bm{\mathrm{f}}}==\text{idx}+1)\cdot\bm{\mathrm{w}}(\text{idx})$
  return $\bm{\mathrm{z}}$
However, this slightly diverts from a “pure” MMV-RAM implementation since it requires to operate on different vectors in parallel. To overcome this, we observe that there actually exists an AC[0] circuit with size $\mathcal{O}(n^{2})$ which performs the REVERT operation, leading to the following result.
Theorem 4.2 (SCD & SSCR).
In MMV-RAM, SCD and SSCR (Algorithms 2 and 3) can be implemented such that they return the segmented sum and scan of $\bm{\mathrm{x}}$ over $\bm{\mathrm{f}}$, respectively, using:
$\mathcal{O}(\log_{s}(n))\text{ steps, }\mathcal{O}\left(\mathcal{M}(\tfrac{n}{s})+nB(n+B)\right)\text{ work,}$
and $B\in\mathcal{O}(\log(nM))$ bits-per-element to avoid overflow, where $M:=\max_{i\in[n]\}}|x(i)|$. Any vector-only MMV-RAM algorithm requires $\Omega\left(\frac{\log(n)}{\log\log(n)}\right)$ steps.
Proof.
[…]
The analysis is almost the same for SSCR (Algorithm 3). The first two steps are identical, since we are performing SCAN and COMPRESS. The only step that differs is the last one, which reverts the speculation. This can be implemented, for example, with an all-to-all communication as in the GATHER step above, with $\mathcal{O}(n^{2}B)$ size and constant depth. ∎
1:procedure SSCR($\bm{\mathrm{x}}$, $\bm{\mathrm{f}}$)
3:  $\bm{\mathrm{f}}^{-}\leftarrow\bm{\mathrm{f}}(1:)$ appended with 11
5:  return $\text{REVERT}(\widehat{\bm{\mathrm{x}}},\widehat{\bm{\mathrm{f}}},\bm{\mathrm{w}})$
Algorithm 3 Segmented Scan $\cong$
REVERT $\circ$ COMPRESS $\circ$ (SCAN, SCAN)
4.2 Improving the work complexity
Theorem 4.2 states that Algorithms 2 and 3 achieve a theoretical speed-up against any VCU-only MMV-RAM algorithm. However, their work scales quadratically with respect to $n$, which is not ideal. Here we show that we can improve the work to $\mathcal{O}(n)$, while maintaining the same step complexity. To achieve this, we need to follow a different approach, which requires specialized circuitry that might not be available on existing hardware. As such, at least at the time of this writing, this improvement is mainly of theoretical interest.
@@ 4.2 Improving the work complexity
In brief, the methodology, which is detailed in Appendix A.2, follows a block-recursive approach to compute segmented scans. The core idea of computing segmented scans (recursively) via a contraction argument is well-known in the literature; see [5] and [53, Algorithms 3 and 4] for circuits with fan-in two. [69] extended the contraction approach to blocks (or fan-in) of size $s$ for unsegmented scan. Our approach is a generalization for the segmented case.
@@
Formally, given an input vector $\bm{\mathrm{x}}$ and a boolean vector $\bm{\mathrm{f}}$ both of size $n$, the block-recursive approach initially partitions $\bm{\mathrm{x}}$ and $\bm{\mathrm{f}}$ into consecutive blocks of size $s$. Then the following steps are executed:
@@
    On each $s$-block of $(\bm{\mathrm{x}},\bm{\mathrm{f}})$, compute (in parallel) the local scans of size $s$ using the MMU.
@@
    Use the VCU (e.g., the dedicated circuit of Lemma A.1) to revert the miss-speculated values, and store the results in place on $\bm{\mathrm{x}}$.
  3. 3.
@@
    Create a vector $\bm{\mathrm{f}}_{s}$ of size $\lceil\tfrac{n}{s}\rceil$ by applying logical OR to the elements of each $s$-block of $\bm{\mathrm{f}}$.
  5. 5.